OpenStudy (anonymous):

What is the distance between points (12, 4) and (18, –1)? Round to the nearest tenth.

OpenStudy (anonymous):

I REALLY NEED THIS ANSWERED PLEASE !

OpenStudy (anonymous):

= sqrt()-1-4)^2 + (18-12)^2) = sqrt (25 + 36) = sqrt 61

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no probs - plz click good answer to close question